OpenStudy (anonymous):

Compare each pair of numbers use <. =. > 3/5 5/8

OpenStudy (anonymous):

3/5 = 24/40 5/8 = 25/40

OpenStudy (anonymous):

3/5 = =(3/5)(8/8)=24/40 5/8 = (5/8)(5/5)=25/40 therefore 25/40 > 24/40 also 5/8 > 3/5 ans.....